OpenStudy (anonymous):

Find sin θ if cot θ = - 4 and cos θ < 0.

OpenStudy (anonymous):

is it -17?

OpenStudy (btaylor):

if cosine and cotangent are negative, it must be in the second quadrant.

OpenStudy (anonymous):

cot(x) - tan(x)/(sin(x) + cos(x) = (cot(x) -tan(x)(sin(x) - cos(x))/(sin^2(x) - cos^2(x) =(cot(x)sin(x) - cot(x)cos(x) - tan(x)sin(x) + tan(x)cos(x))/(sin^2(x) - cos^2(x)) = cos(x) - cos^2(x)/sin(x) - sin^2(x)/cos(x) + sin(x))/(sin^2(x) - cos^2(x)) =( cos^2(x)/cos(x) - sin^2(x)/cos(x) + sin^2(x)/sin(x) - cos^2(x)/sin(x))/(sin^2(x) - cos^2(x)) = ((cos^2(x) - sin^2(x))/cos(x) - (cos^2(x) -sin^2(x))/sin(x)/(sin^2(x) - cos^2(x)) = (cos^2(x) -sin^2(x))(1/cos - 1/sin))/(sin^2(x) - cos^2(x)) = 1/sin(x) - 1/cos(x) = csc(x) - sec(x)

OpenStudy (anonymous):

I guess u r wrong..... @babydoll332

OpenStudy (anonymous):

ohh ok so -1/4

OpenStudy (anonymous):